OpenStudy (anonymous):

Find the maximum and minimum values of x^3 - 6x^2 + 9x + 6 , show carefully how you determine which is which.

OpenStudy (ninjaslice):

This looks better =)

OpenStudy (anonymous):

I think I found the interval for the minimum, 1<x<3, by finding the derivative, using the quadratic formal to find what values make it equal to zero, and using a sign analysis chart.

OpenStudy (anonymous):

If that's correct, I just need help finding the maximum.

OpenStudy (irishboy123):

you mean local maximum, right?  \(y' = 3x^2 - 12x + 9 = 0 \rightarrow x^2 - 4x + 3 =0\rightarrow (x- 1)(x- 3) = 0\)